OpenStudy (anonymous):

Use the rule of 70 to find the approximate doubling time for a population whose annual growth rate is 0.7% (U.S.)? and 1.2% (World)?

OpenStudy (anonymous):

i thought it was the rule of 72! it says doubling time is \[\frac{72}{r}\]

OpenStudy (anonymous):

so first one would be \[\frac{10}{.7}=100\]

OpenStudy (anonymous):

second would be \[\frac{70}{1.2}=58.3\] rounded

OpenStudy (anonymous):

there is rule of 70 I just dont have my book for the formula

OpenStudy (anonymous):

i think 72 is closer, but it is just an approximation in any case

OpenStudy (anonymous):

fine use 70. close enough. it is rule of thumb, not an exact answer.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome!